CALCULATION OF THE MOLAR VOLUME AS A FUNCTION OF PRESSURE FROM THE RAMAN FREQUENCIES IN NH_{4}Br

\underline{H. YURTSEVEN}, Department of Physics, Middle East Technical University, 06531 Ankara, TURKEY.

The molar volume is calculated as a function of pressure using the observed Raman frequencies of the ν_{7} TA (57 cm^{−1}) and ν_{5} TO (140 cm^{−1}) modes of NH_{4}Br in the disordered phase II. This calculation is performed through the mode Grüneisen parameter which is taken as a constant for each mode studied throughout the phase II of this crystal. Measurements of the lattice parameter (molar volume) can be performed to examine our calculations from the Raman frequencies studied here for NH_{4}Br.